# Revisiting Zeroth-Order Optimization for Memory-Efficient LLM Fine-Tuning: A Benchmark

### v2

@@ 4.2 Experiment Results
Inspired by the above, Fig. 3 explores the impact of varying query budget ($q$) on the performance of Forward-Grad and ZO-SGD (i.e., MeZO). As we can see, both the accuracies of Forward-Grad and ZO-SGD are improved with an increased query budget. However, the performance rise for Forward-Grad is much more evident. For example, when the query number is larger than $500$, Forward-Grad outperforms ZO-SGD by a large margin $1\%\sim 2\%$ and approaches FO-SGD. These observations underscore that the inherent advantages of Forward-Grad become apparent only when a sufficient number of queries are used. The downside of using a higher query budget is the increase in computation cost, which scales linearly with $q$.
@@ 5 Extended Study to Improve ZO Fine-Tuning
Block-wise ZO optimization enhances fine-tuning performance.
It has been shown in (Liu et al., 2018) that using a coordinate-wise deterministic gradient estimator can reduce ZO optimization variance, although this scheme is difficult to scale. In a similar vein, we ask if RGE when estimating a FO gradient block-wise can also improve the performance of ZO optimization. The key idea is to split the LLM into different blocks and then apply the ZO gradient estimator to each block of parameters. For example, OPT-1.3B entails $p=26$ parameter blocks, necessitating $p$ forward passes for ZO gradient estimation per fine-tuning step. Our rationale is that by conducting gradient estimation block-wise, the resulting gradient estimate’s variance will be reduced, thereby potentially improving the fine-tuning performance.
@@
Trade-off between performance and memory efficiency via hybrid ZO-FO training. The primary source of memory cost in LLM fine-tuning arises from the BP process, which involves passing gradients from the deep layers to the shallow layers of the model. Hence, to enhance memory efficiency, a potential approach is to confine BP within the deep layers without propagating it to the shallow layers. Moreover, ZO optimization can be employed for the shallow layers without the need for BP. The above approach of combining FO optimization for deep layers and ZO optimization for shallow layers results in a hybrid ZO-FO fine-tuning scheme for LLMs.
@@
Gradient pruning benefits performance.
We next explore gradient pruning, a technique known for accelerating model training without compromising convergence (McDanel et al., 2022). Our key idea is to induce sparse parameter perturbations for reducing gradient estimation variance in RGE. We begin by leveraging magnitude-based pruning (Frankle & Carbin, 2018; Chen et al., 2020) to obtain the layer-wise sparsity ratios. We then generate random pruning masks (following these layer-wise sparsity ratios) and apply them to the weight perturbations in RGE per ZO fine-tuning step. Tab. 7 shows the performance of the sparsity-induced ZO gradient estimation in LLM fine-tuning as a function of the overall sparsity ratio. It becomes evident that choosing a moderate sparsity ratio (e.g., $20\%$) can lead to improved performance over the vanilla ZO optimizer, ZO-SGD.
@@ 6 Conclusion
This work explores the application of zeroth-order (ZO) optimization in fine-tuning large language models (LLMs). ZO optimization approximates gradients using loss differences, eliminating the need for back-propagation and activation storage. While MeZO (Malladi et al., 2023) has made strides in adapting ZO optimization for LLMs, understanding the full ZO landscape remains an open question. To achieve this, we broaden the scope by considering various ZO optimization methods, task types, and evaluation metrics. We conduct the first benchmark study of different ZO optimization techniques, shedding light on their accuracy and efficiency. We also uncover overlooked ZO optimization principles, such as task alignment and the role of forward gradient. Leveraging these insights, we propose techniques like block-wise descent, hybrid ZO and FO training, and gradient sparsity to enhance ZO optimization-based LLM fine-tuning. The proposed enhancements to ZO optimization enable us to further improve fine-tuning accuracy while maintaining memory efficiency.
7 Impact Statements
@@
A Preliminaries of Parameter-Efficient Fine-Tuning (PEFT)
@@
B Zeroth-Order Optimization Algorithms
$\min_{\mathbf{x}\in\mathbb{R}^{n}}f(\mathbf{x})$
^$\hat{\nabla}f(\mathbf{x}):=\frac{f(\mathbf{x}+\mu\mathbf{u})-f(\mathbf{x}-\mu\mathbf{u})}{2\mu}\mathbf{u}$
where $\mathbf{u}$ is a standard Gaussian vector, and $\mu$ is a small scalar. This estimation facilitates the use of gradient-based methods solely based on function evaluations. Utilizing this gradient estimator, we summarize the existing zeroth-other algorithms below.
[…]
$\mathbf{x}_{t+1}=\mathbf{x}_{t}-\eta_{t}\hat{\nabla}f(\mathbf{x}_{t}),$
where $\eta_{t}>0$ is the learning rate.
[…]
$\mathbf{x}_{t+1}=\mathbf{x}_{t}-\eta_{t}\operatorname{sign}(\frac{f(\mathbf{x}+\mu\mathbf{u})-f(\mathbf{x}-\mu\mathbf{u})}{2\mu}\mathbf{u})$
where $\mathbf{u}$ is a standard Gaussian vector.
[…]
$\mathbf{x}_{t+1}=\mathbf{x}_{t}-\eta_{t}\mathbf{m}_{t},$
with the momentum defined as
$\mathbf{m}_{t}=\beta_{t}\mathbf{m}_{t-1}+\hat{\nabla}f(\mathbf{x}_{{t}}).$
ZO-SGD-MMT adopts the momentum factor $\beta_{t}$ to control the average.
[…]
$\mathbf{x}_{t+1}=\operatorname*{arg\,min}_{\mathbf{y}\in\{\mathbf{x}_{t},\mathbf{x}_{t}-\eta_{t}\hat{\nabla}f(\mathbf{x}_{t}),\mathbf{x}_{t}+\eta_{t}\hat{\nabla}f(\mathbf{x}_{t})\}}f(\mathbf{y}).$
[…]
$\mathbf{x}_{t+1}=\mathbf{x}_{t}-\eta_{t}\frac{\sqrt{1-\beta_{2}^{t}}}{1-\beta_{1}^{t}}\mathbf{V}_{t}^{-1/2}\mathbf{m}_{t},$
where the momentum vector ${\mathbf{m}}$ and the second-order momentum matrix $\mathbf{V}$ are computed by
$\mathbf{m}_{t}$ $=\beta_{1}\mathbf{m}_{t-1}+(1-\beta_{1})\hat{\nabla}f(\mathbf{x}_{t}),$
$\mathbf{V}_{t}$ $=\text{Diag}(\mathbf{v}_{t}).$
The hyperparameters $\beta_{1}$ and $\beta_{2}$ are used to control the moving average. Note ZO-Adamm proposed in (Chen et al., 2019) differs from our implementation mainly by the RMSProp, i.e., ${\mathbf{v}}=\max(\mathbf{v}_{t},\mathbf{v}_{t-1})$. We remove RMSProp since it increases the memory cost and could void the advantage of ZO methods compared to FO-SGD.
[…]
$\mathbf{x}_{t+1}=\mathbf{x}_{t}-\eta_{t}(\nabla f(\mathbf{x}_{t})^{\top}\mathbf{u})\mathbf{u},$
where $\mathbf{u}$ is the standard Gaussian random vector. While Forward-Grad utilizes the first-order gradient $\nabla f(\mathbf{x}_{t})$, it employs the Jacobian-Vector Product during the forward pass of training to mitigate computation and memory requirements. As a result, the memory complexity remains relatively low compared to BP-based methods.
@@ C How to Implement Memory-Efficient ZO/FO Optimizers?
  Step 0: Model Loading: Initialize the model with parameter ${\mathbf{x}}$;
  Step 1: Forward Pass: Compute loss $\ell(x)$, and save forward pass states ${\mathbf{s}}_{\text{fwd}}$;
  Step 2: Backward Pass: Calculate gradients w.r.t. ${\mathbf{x}}$, and generate backward states ${\mathbf{s}}_{\text{bwd}}$;
  Step 3: Optimization Step: Update ${\mathbf{x}}$ and ${\mathbf{s}}_{\text{opt}}$ using gradients and utilize temporal state ${\mathbf{s}}_{\text{opt}}^{\prime}$ that will be released immediately;
@@
The memory efficiency of an optimizer heavily depends on its implementation details. This section will discuss these implementation details and provide a holistic memory profile of all the optimizers discussed above. We remark that the discussions in this section are based on the PyTorch framework.
In general, the memory efficiency of an optimizer is defined by the peak memories consumed during training in a specific setting, which primarily depends on the maximum number of variables stored at a certain time point. To dissect the memory consumption of different optimizers, we summarized a general model pipeline for parameter updating in Algorithm A1, which involves four main steps.
First, the program needs to load the model with the full parameters ($\mathbf{x}$) with either full or half-precision. This is an inevitable consumption for all optimizers. Second, a forward pass will yield a loss value and store any involved intermediate states of the model (${\mathbf{s}}_{\text{fwd}}$). Third, the loss in a backward mode utilizes the stored state ${\mathbf{s}}_{\text{fwd}}$. Similarly, this BP process involves storing the backward states ${\mathbf{s}}_{\text{bwd}}$ and temporarily storing $\bm{\tau}_{\text{bwd}}$. After the gradients and state ${\mathbf{s}}_{\text{bwd}}$ are computed, the memory cost from ${\mathbf{s}}_{\text{fwd}}$ will be immediately released. Last, with the computed gradient and states, the parameters and optimizer state will be updated, and the memory of ${\mathbf{s}}_{\text{bwd}}$ will be released. Typically, ${\mathbf{s}}_{\text{bwd}}$ will be the stored gradients. For generality, we assume ${\mathbf{s}}_{\text{fwd}}$ also includes essential temporal variables that are only used in the step.
According to the memory life-cycle, the memory consumption will be summarized as two types: constant memory that exists through the training, and dynamic allocation that only exists in one iteration mainly for gradient computing. Though dynamic allocation only exists temporarily, it also contribute to the peak memory. Without losing generality, we summarize the primary components of the peak memory consumption as

### v3

@@ 4.2 Experiment Results
Table 5: Comparison of the instant peak memory consumption of different optimizers when fine-tuning the full model. Peak memory consumption consists of the constant memory for the model (Weight Mem.), the optimizer states (Opt. State Mem.), and dynamic allocation for computing gradients and optimization (Dynamic Mem.). The cardinalities $|{\mathbf{x}}|$ and $|{\mathbf{a}}|$ correspond to the memory consumption of loading model parameters and saving the intermediate results for post-hoc backward during forward in full precision, with the subscript $l$ representing that of a specific layer $l$. In Dynamic Mem., $\mathbf{x}$ is also used to denote the memory consumption for temporarily saving the gradients, as they share the same size with the corresponding parameters.
Optimizer Weight Mem. Dynamic Mem (Grad.&Opt.) Opt. State Mem.
Optimizer in Full Precision
FO-SGD $|{\mathbf{x}}|$ $\sum_{l}\max\{|{\mathbf{a}}_{l}|,|{\mathbf{x}}_{l}|\}$ 0
FO-Adam w/o fast foreach $|{\mathbf{x}}|$ $\sum_{l}\max\{|{\mathbf{a}}_{l}|,|{\mathbf{x}}_{l}|\}$ $2|{\mathbf{x}}|$
FO-Adam $|{\mathbf{x}}|$ $\sum_{l}\max\{|{\mathbf{a}}_{l}|,|{\mathbf{x}}_{l}|\}$ $3|{\mathbf{x}}|$
Forward-Grad $|{\mathbf{x}}|$ $|{\mathbf{x}}|+\max_{l}|{\mathbf{a}}_{l}|$ 00
Vanilla ZO-SGD $|{\mathbf{x}}|$ $|{\mathbf{x}}|$ 0
ZO-SGD $|{\mathbf{x}}|$ $\max_{l}|{\mathbf{x}}_{l}|$ 0
ZO-SGD MMT $|{\mathbf{x}}|$ $\max_{l}|{\mathbf{x}}_{l}|$ $|{\mathbf{x}}|$
ZO-Adam $|{\mathbf{x}}|$ $\max_{l}|{\mathbf{x}}_{l}|$ $2|{\mathbf{x}}|$
Optimizer in Mixed Precision Training (FP16)
FO-SGD (FP16) $|{\mathbf{x}}|$ $\max\left\{\frac{1}{2}|{\mathbf{a}}|+\frac{1}{2}|{\mathbf{x}}|,\sum_{l}\max\{\frac{1}{2}|{\mathbf{a}}_{l}|,|{\mathbf{x}}_{l}|\}\right\}$ 0
FO-Adam (FP16) $|{\mathbf{x}}|$ $\max\left\{\frac{1}{2}|{\mathbf{a}}|+\frac{1}{2}|{\mathbf{x}}|,\sum_{l}\max\{\frac{1}{2}|{\mathbf{a}}_{l}|,|{\mathbf{x}}_{l}|\}\right\}$ $2|{\mathbf{x}}|$
Optimizer with Half Precision Model (F16)
ZO-SGD (F16) $\frac{1}{2}|{\mathbf{x}}|$ $\max_{l}\frac{1}{2}|{\mathbf{x}}_{l}|$ 0
ZO-SGD-MMT (F16) $\frac{1}{2}|{\mathbf{x}}|$ $\max_{l}\frac{1}{2}|{\mathbf{x}}_{l}|$ $\frac{1}{2}|{\mathbf{x}}|$
ZO-Adam (F16) $\frac{1}{2}|{\mathbf{x}}|$ $\max_{l}\frac{1}{2}|{\mathbf{x}}_{l}|$ $|{\mathbf{x}}|$
4.3 An In-Depth Dissection on Memory Efficiency
In this section, we will provide provide a holistic memory profile for the ZO and FO methods, theoretically and empirically. And we will discuss how memory efficiency will be influenced by the implementation details adopted in this work, including F16 (half-precision model) and FP16 (mixed-precision training).
Theoretical analysis. The theoretical memory efficiency of all the methods with and without the memory-saving tricks are listed in Tab. 5. We refer readers to Appx. C.1 for a detailed analysis of these results. Several key insights can be summarized. First, compared to FO optimizers (including Forward-Grad), the memory efficiency of ZO optimizers mainly comes from two perspectives: (1) ZO methods can avoid saving the intermediate results (model states) $\mathbf{a}_{l}$. As we will see later, this reduction is considerable when the sequence length of the language model is large; (2) ZO methods can estimate the gradient in a layer-wise manner and thus avoid storing the gradients of the entire model. Second, the usage of FP16 in FO methods can only reduce the memory consumption of intermediate results (i.e., $|{\mathbf{a}}|$ in Tab. 5), but not that from the model loading or optimization states’ storage. In contrast, ZO methods can be easily equipped with F16, reducing memory by $50\%$. Third, although Forward-Grad is also a BP-free method, its memory efficiency advantage over other FO optimizers is not remarkable, mainly due to its requirement for storing the intermediate results (i.e., the computational graph). Last, although ZO-Adam is memory intensive compared to ZO-SGD, it can be greatly improved by using F16 and achieves a better efficiency than FO-SGD (FP16).
Empirical results analysis. We further compare the theoretical and empirical memory costs of full fine-tuning in Tab. A1 and LoRA fine-tuning in Tab. A2. Notably, the empirical results are generally aligned with the theoretical analysis. ZO methods are generally much more efficient than their FO counterparts, and the overhead can be further reduced by loading the half-precision model (using F16).
Figure 3: Peak memory comparison of full fine-tuning with FO-SGD and ZO-SGD across various sequence lengths with a fixed effective batch size of $2$. Peak memory consumption was evaluated with the input of synthetic texts generated from random sequences of the specified lengths.
A larger sequence length strengthens the memory benefits of ZO methods. We remark that the empirical results in Tab. A1 and Tab. A2 are dependent on the input sequence length. In general, a larger sequence length will directly lead to a larger activation memory consumption. To investigate its impact, we further explore how the scaling of input sequence length impacts memory usage by increasing activation memory, and the results are shown in Fig. 3. Furthermore, we examine the memory cost of LLM fine-tuning vs. the input sequence length. In Fig. 3, we compare the memory efficiency between ZO-SGD and FO-SGD across various sequence lengths (i.e. the token number per sample). As we can see, the memory consumption of ZO-SGD maintains a consistent level, since its peak memory consumption is only determined by the model parameter size (i.e., the cardinality $|\mathbf{x}|$) and independent of the size of the intermediate results (i.e., $|\mathbf{a}|$), see Tab. 5. In contrast, as the sequence length increases, the peak memory consumption of FO-SGD first maintains and then begins to increase. This phenomenon occurs because the peak memory consumption of FO-SGD is determined by the larger one of the per-layer activation storage and the gradient storage (i.e., $\max\left\{|{\mathbf{a}}_{l}|,|{\mathbf{x}}_{l}|\right\}$). When the sequence length increases, the former begins to grow and eventually exceeds the latter at a certain point (e.g., exceeding 700). Thus, ZO-SGD will exhibit a much better efficiency advantage in settings with long context lengths.
@@ 5 Extended Study to Improve ZO Fine-Tuning
Block-wise ZO optimization enhances fine-tuning performance. It has been shown in (Liu et al., 2018) that using a coordinate-wise deterministic gradient estimator can reduce ZO optimization variance, although this scheme is difficult to scale. In a similar vein, we ask if RGE when estimating a FO gradient block-wise can also improve the performance of ZO optimization. The key idea is to split the LLM into different blocks and then apply the ZO gradient estimator to each block of parameters. For example, OPT-1.3B entails $p=26$ parameter blocks, necessitating $p$ forward passes for ZO gradient estimation per fine-tuning step. Our rationale is that by conducting gradient estimation block-wise, the resulting gradient estimate’s variance will be reduced, thereby potentially improving the fine-tuning performance.
@@
Trade-off between performance and memory efficiency via hybrid ZO-FO training. The primary source of memory cost in LLM fine-tuning arises from BP, which involves passing gradients from the deep layers to the shallow layers of the model. To save memory, a potential approach is to confine BP within the deep layers without propagating it to the shallow layers. Moreover, ZO optimization can be employed for the shallow layers without the need for BP. The above approach of combining FO optimization for deep layers and ZO optimization for shallow layers results in a hybrid ZO-FO fine-tuning scheme for LLMs.
@@
Gradient pruning benefits performance. We next explore gradient pruning, a technique known for accelerating model training without compromising convergence (McDanel et al., 2022). Our key idea is to induce sparse parameter perturbations for reducing gradient estimation variance in RGE. We begin by leveraging magnitude-based pruning (Frankle & Carbin, 2018; Chen et al., 2020) to obtain the layer-wise sparsity ratios. We then generate random pruning masks (following these layer-wise sparsity ratios) and apply them to the weight perturbations in RGE per ZO fine-tuning step. Tab. 8 shows the performance of the sparsity-induced ZO gradient estimation in LLM fine-tuning as a function of the overall sparsity ratio. It becomes evident that choosing a moderate sparsity ratio (e.g., $20\%$) can lead to improved performance over the vanilla ZO optimizer, ZO-SGD.
@@ 6 Conclusion
This work explores the application of zeroth-order (ZO) optimization in fine-tuning LLMs. ZO optimization approximates gradients using loss differences, eliminating the need for back-propagation and activation storage. While MeZO (Malladi et al., 2023) has made strides in adapting ZO optimization for LLMs, understanding the full ZO landscape remains an open question. To address this question, we broaden the scope by considering various ZO optimization methods, task types, and evaluation metrics. We conduct the first benchmark study of different ZO optimization techniques, shedding light on their accuracy and efficiency. We also uncover the overlooked ZO optimization principles, such as task alignment and the role of forward gradient. Leveraging these insights, we propose techniques like block-wise descent, hybrid ZO and FO training, and gradient sparsity to enhance ZO optimization-based LLM fine-tuning. The proposed enhancements can further improve the fine-tuning accuracy while maintaining the memory efficiency.
Impact Statement
@@
Acknowledgement
The work of Y. Zhang and S. Liu is partially supported by the U.S. Department of Energy via Lawrence Livermore National Laboratory. The work of J. D. Lee and Z. Wang is in part supported by the NSF AI Institute for Foundations of Machine Learning (IFML). The work of M. Hong and J. Li is partially supported by NSF grant CCF-1910385 and CNS-2003033. This work is also supported by an Amazon AWS gift award.
@@ Acknowledgement
A Zeroth-Order Optimization Algorithms
$\min_{\mathbf{x}\in\mathbb{R}^{n}}f(\mathbf{x})$
^$\hat{\nabla}f(\mathbf{x}):=\frac{f(\mathbf{x}+\mu\mathbf{u})-f(\mathbf{x}-\mu\mathbf{u})}{2\mu}\cdot\mathbf{u}$
where $\mathbf{u}$ is a standard Gaussian vector, and $\mu$ is a small scalar. This estimation facilitates the use of gradient-based methods solely based on function evaluations. Utilizing this gradient estimator, we summarize the existing zeroth-other algorithms as follows
[…]
$\mathbf{x}_{t+1}=\mathbf{x}_{t}-\eta_{t}\hat{\nabla}f(\mathbf{x}_{t}).$
[…]
$\mathbf{x}_{t+1}=\mathbf{x}_{t}-\eta_{t}\operatorname{sign}(\frac{f(\mathbf{x}+\mu\mathbf{u})-f(\mathbf{x}-\mu\mathbf{u})}{2\mu})\mathbf{u}$
where $u$ is a standard Gaussian vector.
[…]
$\mathbf{x}_{t+1}=\mathbf{x}_{t}-\eta_{t}\mathbf{m}_{t},$
with the momentum defined as
$\mathbf{m}_{t}=\beta_{t}\mathbf{m}_{t-1}+\hat{\nabla}f(\mathbf{x}_{{t}}).$
[…]
$\mathbf{x}_{t+1}=\operatorname*{arg\,min}_{\mathbf{y}\in\{\mathbf{x}_{t},\mathbf{x}_{t}-\eta_{t}\hat{\nabla}f(\mathbf{x}_{t}),\mathbf{x}_{t}+\eta_{t}\hat{\nabla}f(\mathbf{x}_{t})\}}f(\mathbf{y})$
Due to the large variance introduced by the zeroth-order RGE estimator, we believe that the conservative update could correct the wrong step made by ZO-SGD, yielding better convergence results.
[…]
$\mathbf{x}_{t+1}=\mathbf{x}_{t}-\eta_{t}\mathbf{V}_{t}^{-1/2}\mathbf{m}_{t},$
where the momentum $\mathbf{m}_{t}$, second raw momentum estimate $\mathbf{v}$ and normalization matrix $\mathbf{V}$ are defined as
$\mathbf{m}_{t}$ $=\beta_{1,t}\mathbf{m}_{t-1}+(1-\beta_{1,t})\hat{\nabla}f(\mathbf{x}_{t}),$
$\mathbf{V}_{t}$ $=\text{Diag}(\max(\mathbf{v}_{t},\mathbf{v}_{t-1})).$
[…]
$\mathbf{x}_{t+1}=\mathbf{x}_{t}-\eta_{t}(\nabla f(\mathbf{x}_{t})^{\top}\mathbf{u})\mathbf{u}$
, where $\mathbf{u}$ is the standard Gaussian random vector. Though the Forward Gradient uses the first-order gradient $\nabla f(\mathbf{x}_{t})$, it uses the Jacobian-Vector Product (JVP) during the forward pass of training to reduce the computation and memory consumption. Therefore, the memory complexity is still comparably low compared to back-propagation-based approaches.
B Preliminaries of Parameter-Efficient Fine-Tuning (PEFT)
@@ C How to Implement Memory-Efficient ZO/FO Optimizers?
  Step 0: Model Loading
  $\bullet$ Function: Initialize the model with parameter ${\mathbf{x}}$;
  $\bullet$ Memory Consumption: ${\mathbf{x}}$;
  $\bullet$ Memory Release: N/A.
  Step 1: Forward Pass
  $\bullet$ Function: Compute loss $\ell(x)$, save forward pass states ${\mathbf{s}}_{\text{fwd}}$, and use temporary variables $\bm{\tau}_{\text{fwd}}$.
  $\bullet$ Memory Consumption: $\ell(x),{\mathbf{s}}_{\text{fwd}},\bm{\tau}_{\text{fwd}}$;
  $\bullet$ Memory Release: Release $\bm{\tau}_{\text{fwd}}$ after computation.
  Step 2: Backward Pass
  $\bullet$ Function: Calculate gradients w.r.t. ${\mathbf{x}}$, generate backward states ${\mathbf{s}}_{\text{bwd}}$, employ temporary variables $\bm{\tau}_{\text{bwd}}$;
  $\bullet$ Memory Consumption: ${\mathbf{s}}_{\text{bwd}},\bm{\tau}_{\text{bwd}}$;
  $\bullet$ Memory Release: Release ${\mathbf{s}}_{\text{fwd}},\bm{\tau}_{\text{bwd}}$ after gradients are computed.
  Step 3: Optimization Step
  $\bullet$ Function: Update ${\mathbf{x}}$ and ${\mathbf{s}}_{\text{opt}}$ using gradients, utilize temporary variables $\bm{\tau}_{\text{opt}}$.
  $\bullet$ Memory Consumption: ${\mathbf{x}},{\mathbf{s}}_{\text{opt}},\bm{\tau}_{\text{opt}}$.
  $\bullet$ Memory Release: Release ${\mathbf{s}}_{\text{bwd}},\bm{\tau}_{\text{opt}}$ post-update.
@@
The memory efficiency of an optimizer heavily depends on its implementation details. In this section, we will discuss these implementation details and provide an holistic memory profile of all the optimizers discussed above. We remark that the discussions in this section are based on the PyTorch framework. In general, the memory efficiency of an optimizer is defined by the peak memories consumed during training in a specific setting, which primarily depends on the maximum number of variables stored at a certain time point. To dissect the memory consumption of different optimizers, we summarized a general model parameter updating pipeline in Algorithm A1, which involves four main steps. First, the program needs to load the model with the full parameter ${\mathbf{x}}$ with either full or half-precision. This is an inevitble consumption for all optimizers. Second, a forward pass will yield a loss value and store any involved intermediate state of the model ${\mathbf{s}}_{\text{fwd}}$. This process may also requires storing any other temporary variables $\bm{\tau}_{\text{fwd}}$ (e.g. an extra copy in float16 of model weight in FO-SGD FP16), which will be immediately released after the forward is complete. Third, the loss is back-propagated in a backward mode utilizing the stored state ${\mathbf{s}}_{\text{fwd}}$. Similarly, this process involves storing the backward states ${\mathbf{s}}_{\text{bwd}}$ and temporarily storing $\bm{\tau}_{\text{bwd}}$. After the gradients and state ${\mathbf{s}}_{\text{bwd}}$ are computed, the memory cost from ${\mathbf{s}}_{\text{fwd}}$ will be immediately released. Last, with the computed gradient and states, the parameters and optimizer state will be updated, and the memory of ${\mathbf{s}}_{\text{bwd}}$ will be released. Typically, ${\mathbf{s}}_{\text{bwd}}$ will be the stored gradients. Since the forward states and gradients are computed layer-wise, the memory consumption will be sequentially transmitted from the state to the gradient. According to the life-cycle, the memory consumption will be summarized as two types: constant memory that exists through the training life-cycle, and dynamic allocation that only exists in one iteration mainly for gradient computing. Though dynamic allocation only temporarily exist, it can increase the peak memory. As a result, we summarize the peak memory consumption as